# Nocturnal Eye-Inspired Liquid-to-Gas Phase Change Soft Actuator with Laser-Induced Graphene: Enhanced Environmental Light Harvesting and Photothermal Conversion

### v2

@@ 4.1 Experimental analysis of actuation driven by light On-Off cycling
The experimental setup is illustrated in Fig. 5. During the experiments, the fabricated actuator was positioned 50 mm away from the light source. The light bulb specifications include luminous intensity of 300 cd, total luminous flux of 630 lm, and beam angle of $60^{\circ}$. An iPhone 15 Pro Max (Apple Inc., Cupertino, CA, USA) was used to capture the actuator deformation, and ImageJ (citation) was employed for analyzing the bending angles and deformation measurements.
@@
Figure 6A and C show the bending behavior during light irradiation, the subsequent recovery upon light off, and the corresponding analysis results. When exposed to light, the without-LIG actuator exhibited delayed activation, with a 63% response time of 141.7 seconds and approximately 220 seconds required to reach the maximum bending angle. In contrast, the proposed actuator with LIG demonstrated a mean 63% response time of 65.0 seconds. It reached maximum bending angle in approximately 90 seconds, showing a 54.1% improvement in responsiveness compared to the without-LIG actuator.
@@
Regarding the recovery phase after light off (Fig 6B and D, the without-LIG actuator showed a 63% response time of 42.7 seconds and required 120 seconds to return to its initial state. The proposed actuator with LIG demonstrated enhanced recovery characteristics with a 63% response time of 22.1 seconds and approximately 60 seconds to return to its initial state, representing a 48.2% improvement in recovery performance compared to the without-LIG actuator.
4.2 Deformation of the actuator before and after light-driven operation
@@
Figure 7 demonstrates the variations in actuator performance based on changes in the encapsulated liquid volume and the distance between the light source and actuator. With the light source fixed at 50 mm, we examined the effect of varying volumes (20 µL, 30 µL, and 40  µL) of encapsulated Opteon SF33. The results showed volume-dependent increases in three parameters: the rate of actuator deformation (Fig. 7A), the plateau bending angle (Fig. 7B), and the time required to reach the plateau state (Fig. 7C). The linear regression analysis yielded an R² value of approximately 0.98, indicating that the relationship can be effectively approximated as linear within the 20-40 µL range.
The influence of external environmental conditions was investigated by varying the distance between the light source and the actuator (50 mm, 75 mm, and 10 mm), with results presented in Fig. 7D-F (Opteon SF33’s amount is 30 µL). The deformation ratio and plateau bending angle decreased as the distance increased. Notably, the time required to reach the plateau state was the longest in the 100 mm group, showing more than a 2.5-fold increase compared to the 50 mm group.
@@ 5 Multi-Finger Robot Using Non-Contact & Light-Driven LIG-Transferred Soft Actuators
To verify the synchronous actuation capability of multiple LIG-transferred soft actuators, we developed a three-fingered robot inspired by the Venus flytrap and evaluated its operation in ambient conditions. The demonstration is shown in Movie S2 and Fig. 8. As illustrated in Fig. 8, upon irradiation from a light source (a lamp in this experiment), the actuators began to bend synchronously, successfully enveloping a ladybug-inspired target object through their coordinated curving motion. When the light source was switched off, each actuator was released from its bent state and returned to its original configuration.
@@ 6.2 Gas leakage
As shown in Fig. 7F, increasing the distance from the light source decreased the bending angle at a plateau. This suggests that Opteon SF33 is leaking out due to silicone’s high gas permeability during the slower actuation caused by reduced light intensity. Similar gas leakage phenomena have been discussed in previous studies[15]. The current actuator design implements a reusable mechanism where one side is sealed with a clip, allowing multiple liquid refilling. However, frequent leakage can limit potential applications. Potential approaches include incorporating materials with lower gas permeability and developing alternative sealing mechanisms to address this limitation.
@@ 6.3 Utilization of ambient environmental light
The light source utilized in this study was a reflector lamp (54W) with a beam angle of $60^{\circ}$, center luminous flux of 630 lm, and beam flux of 160 lm. Assuming uniform light distribution, calculations based on illuminance indicate that the actuator surface, positioned at a distance of 50 mm, receives approximately 60,000 lux (assuming an actuator LIG surface area of 40 $\rm{mm}^{2}$). This illuminance corresponds to approximately 60% of direct summer sunlight[29]. While outdoor operation would depend on seasonal and geographical factors, implementing lens-based light concentration mechanisms could make outdoor applications feasible. Given that typical office environments maintain illuminance levels of approximately 300–1000 lux[30], indoor operation under ambient lighting conditions would be significantly limited. Future investigations will explore localized light energy irradiation using laser sources aimed at achieving more precise actuator control at a microscale level.
@@ 7 Conclusions
This paper focused on the ocular structure of nocturnal animals that efficiently collect and utilize light in low-irradiation environments. We addressed the limitation of conventional silicone-based soft actuators, which exhibited low driving efficiency due to ineffective light utilization. We incorporated a light-heat conversion element layer called LIG on the light-incident surface. This modification enabled the actuator to effectively utilize light that would otherwise be poorly absorbed by silicone alone, heating the internal low-boiling-point liquid. As a result, the proposed actuator with LIG demonstrated a 54% improvement in response time compared to conventional without-LIG actuators.
While various approaches have successfully enhanced photothermal conversion and thermal conductivity by incorporating graphene or metal powders into silicone elastomers, our method explores an alternative approach by installing a LIG layer on the silicone surface. This approach preserves the inherent properties of silicone while utilizing LIG, which can be manufactured under standard environmental conditions, thereby facilitating both the production of photothermal actuators and their diverse applications.
Future research will focus on the versatility of LIG transfer patterns, aiming to develop efficient light-utilizing LIG patterns under natural environmental conditions and design more complex actuator geometries with corresponding LIG patterns. This research direction aligns with the principles of soft robotics, pursuing the development of soft actuators with enhanced deformation capabilities and environmental adaptability.
@@ Acknowledgement
This work was supported by JSPS KAKENHI Grant Number JP24K21213. M.S. sincerely thank my cats, whose eyes served as a visual reference for the demonstration of the tapetum lucidum structure and whose presence inspired and supported me throughout the research process.
@@ 8 Supplementary Material
    Supplementary Figure S1
@@
    Supplementary Movie The actuator’s bending under light irradiation and its bending release upon turning off the light.
@@
    Supplementary Movie S2:Demonstration of a bio-inspired (Venus flytrap) robot utilizing the proposed actuator in a biomimetic movement driven by light.

### v4

@@ 4.1 Experimental analysis of actuation driven by light On-Off cycling
The experimental setup is illustrated in Fig. 5. During experiments, the fabricated actuator was positioned 50 mm from the light source. The light bulb had a luminous intensity of 300 cd, a total luminous flux of 630 lm, and a beam angle of 60°. An iPhone 15 Pro Max (Apple Inc., Cupertino, CA, USA) captured actuator deformation, and ImageJ analyzed bending angles and deformation measurements[36].
@@
Figure 6A, C and Supplementary Movie S1 show bending behavior during light irradiation, subsequent recovery upon light off, and corresponding analysis results. When exposed to light, the without-LIG actuator exhibited a 63% response time of 141.7 seconds and required approximately 220 seconds to reach maximum bending. In contrast, the proposed actuator with LIG demonstrated a mean 63% response time of 65.0 seconds and reached maximum bending in about 90 seconds, a 54.1% improvement.
@@
During recovery after light off (Fig. 6, D and Supplementary Movie S1), the without-LIG actuator showed a 63% response time of 42.7 seconds and needed 120 seconds to return to its initial state. The proposed actuator with LIG demonstrated a 63% response time of 22.1 seconds and took about 60 seconds to return, a 48.2% improvement in recovery performance.
To assess repeatability under cyclic light conditions, we performed multiple actuation tests with a single fluid charge (Fig. 7). We encapsulated 40 µL of Opteon SF33 and exposed the actuator to repeated light irradiation. The bending angles during the first three cycles were 51.7°, 51.6°, and 50.7°, demonstrating good consistency. In the fourth cycle, however, the bending angle declined to 90.4% of the initial peak value and, by the sixth cycle, decreased to 50.5%. We attribute this degradation to gas leakage through the silicone elastomer’s gas permeability. When we immersed the actuator in water at 65C,, gas bubbles emerged from expanded and thinned areas of the silicone wall (Supplementary Figure 2). To address this limitation, we designed the actuator with a refillable configuration and incorporated a detachable clip-based sealing mechanism for repeated operation.
4.2 Analysis of Actuator Performance Under Different Liquid Volume and Light Distance Conditions
@@
Figure 8 demonstrates the variations in actuator performance based on changes in the encapsulated liquid volume and the distance between the light source and actuator. With the light source fixed at 50 mm, we examined the effect of varying volumes (20 µL, 30 µL, and 40 µL) of encapsulated Opteon SF33. The results showed volume-dependent increases in three parameters: the angular change ratio (Fig. 8A), the plateau bending angle (Fig. 8B), and the time required to reach the plateau state (Fig. 8C). The linear regression analysis yielded an $R^{2}$ value of approximately 0.98, indicating that the relationship can be effectively approximated as linear within the 20–40 µL range.
We examined how environmental conditions affect the actuator by testing it at different distances from the light source 50 mm, 75 mm, and 100 mm. The results are shown in Fig. 8D–F using 30µL of Opteon SF33. As we moved the actuator further from the light source, both the angular change ratio and the plateau bending angle became smaller. The actuator positioned at 100 mm from the light source required a significantly longer time to reach its final bent configuration, approximately 2.5 times greater than that observed at 50 mm distance.
5 Mechanical characteristics of the actuator
We evaluated the load-bearing performance of the actuator by attaching polyurethane sponge weights to its distal tip and measuring the resulting reduction in bending angle(Fig. 9A). A nylon thread was connected to the tip of a pre-bent actuator, and sponge-based weights of 0.125, 0.25, 0.5, 0.75, 1.0, and 1.5 g were sequentially suspended via a fine pin. The results (Fig. 9B) indicate that the actuator maintained approximately 90% of its original bending angle under a 0.125 g load ($\approx$ 1.23 mN). As the load increased, the actuator’s ability to retain its bent configuration declined, reaching around 50% of the original angle under a 1.5 g load ($\approx$ 14.7 mN). These results demonstrate a near-linear relationship between the applied load and bending angle reduction, providing a quantitative evaluation of the actuator’s mechanical output capacity.
@@ 6 Multi-Finger Robot Using Non-Contact & Light-Driven LIG-Transferred Soft Actuators
Investigating the practical applications of this actuator, we conducted experimental validation of the multi-finger actuator’s performance in controlled laboratory conditions and assessed its functionality under natural solar illumination.
6.1 Three-finger actuator performance in laboratory conditions
To evaluate the feasibility of the proposed actuator system, we investigated synchronized operation of multiple actuators under controlled light exposure conditions and examined the effects of photoactivation and deactivation on actuator performance. We assessed the coordinated bending motion of three actuators in response to a light stimulus, aiming to achieve simultaneous coverage of a designated target object. The experimental results are documented in Supplementary Movie S2 and Supplementary Figure S3. The experimental setup comprised a controlled laboratory environment. Illumination was provided by reflector lamps throughout experimental trials.
The experimental findings demonstrated that upon exposure to incident light at an intensity of approximately 15 klx, all three actuators exhibited synchronized bending behavior with high temporal coordination. Upon cessation of light exposure, the actuators consistently returned to their initial linear configuration, confirming the reversible nature of the photoresponsive actuation mechanism.
6.2 Solar-responsive shading device demonstration
In natural environments, intense solar radiation can cause harm to flora and fauna. To address this challenge, we developed a prototype shading device (Fig. 10) that responds directly to solar irradiance by providing shade coverage over designated target organisms. The device architecture incorporates three actuators integrated with nonwoven fabric. We deployed this system outdoors to evaluate solar-driven performance. Testing was conducted at 14:00 JST under clear atmospheric conditions. On the testing date, the ambient illuminance measured 102.5 klx. Using clover as the target organism, we positioned the device under direct solar exposure. As demonstrated in Fig. 10 and Supplementary Movie S3, the actuators responded to sunlight by bending and successfully providing shade coverage over the clover. Following deployment, illuminance within the shaded region decreased to 37.6 klx, achieving a 63.3% reduction. The actuation process required approximately 140 seconds to complete.
@@ 7.2 Gas leakage
As shown in Supplementary Fig. 2, gas leakage phenomena were confirmed to occur in the proposed actuator, similar to previously reported issues[15]. The current actuator design features a reusable mechanism with one side sealed by a clip, allowing for multiple liquid refilling cycles. Future work requires optimal design of the refilling port to enhance sealing performance and improve airtightness. However, gas leakage can limit potential applications. Potential approaches include incorporating materials with lower gas permeability and developing alternative sealing mechanisms to address this limitation.
@@ 7.3 Utilization of ambient environmental light
The optical characterization and performance evaluation were conducted under controlled laboratory conditions and outdoor environments. In the laboratory setup, a reflector lamp (54 W, beam angle: 60°) was employed as the primary light source, exhibiting a center luminous flux of 630 lm and a beam luminous flux of 160 lm. Illuminance measurements yielded approximately 15 klx at a working distance of 100 mm and 30 klx at 50 mm, respectively.
Comparative outdoor experiments under direct solar irradiation demonstrated higher illuminance levels of 100–110 klx, indicating laboratory conditions represented about 30% of natural outdoor intensity. The actuator demonstrated slower bending speed outdoors despite increased light intensity, requiring approximately 140 seconds to reach maximum bending angle (Fig. 10). This slower performance stems from convective heat transfer caused by wind, which cools the actuator and prevents the necessary temperature rise. Outdoor deployment faces challenges from seasonal changes, location factors, and weather conditions, limiting fast response times. Optical concentration systems, such as focusing lenses, could boost light energy density on the LIG surface. Smart use of wind-induced cooling might enable quicker bending–unbending cycles, broadening outdoor applicability.
Additionally, localized light energy irradiation using laser sources to improve actuator control precision at the microscale represents a promising avenue for future research.
@@ 8 Conclusions
We addressed the limitation of conventional silicone-based soft actuators, which exhibited low driving efficiency due to ineffective light utilization. We incorporated a photothermal conversion layer using LIG on the light-incident surface. This modification enabled the actuator to effectively utilize light that would otherwise be poorly absorbed by silicone alone, heating the internal low-boiling-point liquid. As a result, the proposed actuator with LIG demonstrated a 54% improvement in response time compared to conventional without-LIG actuators.
Our LIG-based approach preserves silicone’s inherent characteristics while enabling fabrication under ambient conditions and operation with environmental light sources including sunlight. This method significantly simplifies the manufacturing process of photothermal actuators and opens up promising possibilities for their broader application in diverse fields.
@@ Acknowledgement
This work was supported by the Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Number JP24K21213. This work was supported by "Advanced Research Infrastructure for Materials and Nanotechnology in Japan(ARIM)" of the Ministry of Education, Culture, Sports, Science and Technology (MEXT), Grant Number JPMXP1225UT0165. M.S. sincerely thanks T. Kondo, who kindly provided support in the use of the scanning electron microscope, and my cats, whose eyes served as a visual reference for demonstrating the tapetum lucidum structure and whose presence inspired and supported me throughout the research process.
@@ 9 Supplementary Material
    Supplementary Figure S1:Effect of light irradiation direction Temperature changes on the opposite surface (liquid-contact surface) when the light was irradiated from one direction onto either the silicone side (A) or LIG side (B) of silicone elastomer with transferred LIG or silicone only (C). The fastest temperature rise was observed when light was irradiated from the silicone side (case A).
@@
    Supplementary Figure S2: Visualization of gas leakage from the proposed actuator. The proposed actuator was immersed in 65C hot water to visualize gas leakage from within. Small bubbles can be seen leaking from the expanded actuator.
@@
    Supplementary Figure S3: Synchronized actuation of three-finger actuator system under light control. A: Schematic illustration showing the operating principle: actuators remain straight when light is OFF and bend to form a protective enclosure when light is ON through a "hold and protect" mechanism. B: Time-lapse demonstration of synchronized actuation in laboratory conditions showing three actuators simultaneously bending to cover a target object upon light exposure (15 klx). At t = 0.00 seconds (light-off), the actuators maintain a straight configuration. At t = 1.00 seconds (light-on), the actuators exhibit a coordinated bending motion, forming a protective enclosure. At t = 2.20 seconds (light-off), the actuators return to their initial straight state, demonstrating reversible and controllable actuation through light activation and deactivation.
  4. 4.
    Supplementary Movie S1: The actuator’s bending under light irradiation and its bending release upon turning off the light (related to Fig. 6).
  5. 5.
    Supplementary Movie S2: Demonstration of a multi-finger robot in a laboratory setup mimicking a heterogeneous natural environment. The ambient temperature was 20C, and the illuminance was 30.2 klx. Under light irradiation, the three actuators moved synchronously, and upon light-off, their bending was released (see Supplementary Figure S3).
  6. 6.
    Supplementary Movie S3: Demonstration of shading robot driven by sunlight in nature (related to Fig. 10).